OpenStudy (anonymous):

Find two integers whose sum is 26 and whose product is maximum.

OpenStudy (anonymous):

make them equal by symmetry

OpenStudy (anonymous):

What do do you mean?

OpenStudy (anonymous):

suppose i call one integer \(x\) and the other integer \(y\) then you have \[x+y=26\] or \[y=36-x\] and you want to maximize \[xy=x(26-x)\]

OpenStudy (anonymous):

typo there, i meant \(y=26-x\) but you get the idea

OpenStudy (anonymous):

it is easy enough to use algebra to maximize \[A(x)=x(26-x)=26x-x^2\] to find the vertex of the parabola but it is easier to think as follows

OpenStudy (anonymous):

I think I understand. So what does the maximum mean in the problem?

OpenStudy (anonymous):

find the value of \(x\) that makes \[26x-x^2\] as large as possible i.e. find the vertex of the parabola

OpenStudy (anonymous):

but as i said, it is a lot easier than doing that since you cannot tell x from y in this question, you get the same answer if you replace y by x therefore they must be equal, i.e. they are both 13